I have tried over and over and still can not get this problem right. I am trying to do a linear inequalities by graphing.
x-3y >6
3x+2y >12

Answer by benni1013(206)

Inequalities are treated just like equations except for one thing which I will tell you about later.
x-3y>6
3x+2y>12
Turn the inequalities into equations.
x-3y=6
3x+2y=12
Turn the equations from standard form into slope intercept form
x-3y=6==>-3y=6-x==>y=-2+1/3x==>y=1/3x+2
3x+2y=12==>2y=12-3x==>y=6-3/2x==>y=-3/2x+6
Graph the point (0,2) and go up one and over three and plot that point.
Now the difference between graphing inequalities and equations.
When you have either the greater than or less than sign use a dashed line. If you have either a greater than or equal to or a less than or equal to sign use a solid line. If the sign is greater than shade to the right if it is less than shade to the left. In both cases you will draw a dashed line and draw it to the right.